OpenStudy (anonymous):

Simplify the given expression. (7 − 3i) • (2 − i) 4 + i 14 + 3i 11 − 13i 14 − 10i

OpenStudy (anonymous):

So... We distribute so that... 7 * 2 = 14 7 * i = 7i then -3i * 2 = -6i -3i * -i = 3i?

OpenStudy (anonymous):

then combine like terms... 7i - 6i + 3i = 4i so 14 + 4i?

OpenStudy (anonymous):

dang that isn't a possible answer :S

terenzreignz (terenzreignz):

You MUST recall that \[\Large i^2 = -1\]

OpenStudy (anonymous):

wait a minute! -3i * i = 3i^2 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

7i - 6i + 3i^2... 1i + 3i^2 is idk...

OpenStudy (anonymous):

and i^2 = -1, so you have 3(-1)-13i+14 = -3-13i+14=11-13i

OpenStudy (anonymous):

3(-1) -3 so 1 -3 -2?

OpenStudy (anonymous):

lol I can't get this xD

OpenStudy (anonymous):

you're getting your negative numbers wrong.....-7 times i is -7i

OpenStudy (anonymous):

hmm... 7 * 2 = 14 7 * i = 7i then -3i * 2 = -6i -3i * -i = 3i^2 3i^2 is 3(-1) so -3... 14 - 3 11 then 7i - 6 = 1i or i

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

7 * -i = -7i

OpenStudy (anonymous):

11 - 13i?

OpenStudy (anonymous):

did i get it :D?

OpenStudy (anonymous):

FOIL First, inside, outside, last (7)(2)=14 (-3i)(2)=-6i (7)(-i)=-7i (-3i)(-i)=3i^2 Combine like terms 3i^2-13i+14 i^2 in imaginary numbers=-1 3(-1)-13i+14=-13i+11 or 11-13i

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

Thanks for the help man

OpenStudy (anonymous):

You're welcome!